OpenStudy (anonymous):

find the slope of the line trough each pair of points (4, -1) and (-2, -3)

OpenStudy (anonymous):

\[m=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

Yea that didnt help at all

OpenStudy (anonymous):

 slope = m = tan(theta) = b/a = 2/6= 1/3